OpenStudy (anonymous):

Isabel is running. The number of calories she has burned varies directly with the number of minutes she has run. in 10 min 100 calories burned in 20 min 200 calories burned in 30 min 300 calories burned

OpenStudy (anonymous):

how many min did isbel run per calories burned

OpenStudy (kirbykirby):

\[\frac{10~min}{100 ~calories}=\frac{x~min}{1~calorie}\]

OpenStudy (anonymous):

A freight train is carrying goods across the country. The number of gallons of fuel it has used varies directly with the distance it has traveled.

OpenStudy (anonymous):

30 gallon at 50 mile 60 gallon at 100 mile 90 gallon at 150 mile how many gallons does the train use per mile

OpenStudy (anonymous):

@kirbykirby can you help

OpenStudy (kirbykirby):

it's a similar question to the previous one. Notice that doubling the gallons also doubles the miles (from 30 to 60 gallons)? Everything is directly proportional. So \[\frac{30~gallons}{50~miles}=\frac{x~gallons}{1~ mile} \]

OpenStudy (kirbykirby):

If they ask what you get PER (unit), like here it's PER mile, then this just means "one", like per mile means, for every "one" mile